A body moving in a progressive wave pattern has acceleration of 10cm/s² when the displacement is 6cm and a speed of 12cm/s² when the displacement is 8cm. What is the amplitude of the wave?
1
Expert's answer
2019-11-27T10:18:30-0500

Amplitude of the Wave

We need to find the Amplitude of the wave.


Solution:


Given,

Acceleration = "a =10 \space cm/s²"


Displacement = "x = 6 \space cm"


Frequency = "\omega"


We know the formula,

"a = \omega^2 \times x"

Plug the values in this,


"10 = \omega ^2 \times 6"


"\omega^2 = \frac {10}{6}"

"\omega = \sqrt {\frac {10} {6} }"

Here,

"Speed = v = 12 \space cm/sec \\
Displacement = x = 8 \space cm"

"v = \omega \space \sqrt { A^2 - x^2}"

Squaring on both sides,

"v^2 = \omega ^2 ( A^2 - x^2)"

Here, A is the Amplitude,


Plug the values in the formula,


"144 = (\frac {10}{6}) ( A^2 - 64)"

"A^2 - 64 = 144 \times \frac {6}{10} = 86.4"

"A^2 = 86.4 + 64 = 150.4"

So,

"A = \sqrt {150.4} = 12.26 \space cm"

Answer: Amplitude of the Wave = A = 12.26 cm